# Communication-Efficient Decentralized Optimization via Double-Communication Symmetric ADMM

### v1

@@ Abstract
This paper focuses on decentralized composite optimization over networks without a central coordinator. We propose a novel decentralized Symmetric ADMM algorithm that incorporates multiple communication rounds within each iteration, derived from a new constraint formulation that enables information exchange beyond immediate neighbors. While increasing per-iteration communication, our approach significantly reduces the total number of iterations and overall communication cost. We further design optimal communication rules that minimize the number of rounds and variables transmitted per iteration. The proposed algorithms are shown to achieve linear convergence under standard assumptions. Extensive experiments on regression and classification tasks validate the theoretical results and demonstrate superior performance compared to existing decentralized optimization methods. To our knowledge, this is the first decentralized optimization framework that achieves a net reduction in total communication by leveraging fixed multi-round communication within each iteration.
@@ 1 Introduction
In decentralized optimization, a central challenge lies in reducing the time cost of both local computation and inter-node communication. While existing algorithms differ in their local update strategies, most follow a common structural pattern: each iteration is followed by a single round of communication. This convention has seldom been challenged, primarily due to the concern that introducing multiple communication rounds per iteration would increase the overall communicational cost. Prior attempts to incorporate fixed multiple communication rounds, such as those in (Multi1, ; DGDT, ; ProxMudag, ; Multi2, ), achieve this by multi-consensus—repeatedly mixing local variables through communication. However, these methods have not demonstrated practical reductions in the total number of communication rounds. As a result, the potential for achieving a net communication reduction through non-adaptive multi-communication algorithms remains largely unexplored.
@@
In this paper, we investigate the potential of integrating multiple rounds of communication within a single iteration motivated by this observation. Rather than directly applying multiple mixing steps, we develop our algorithms by introducing linear constraints tailored for ADMM, which naturally embed multi-round communication into each iteration. To further enhance performance, we adopt a Symmetric ADMM (SADMM, ) framework to accelerate convergence. Although this design increases the per-iteration communication cost, it enables a significantly faster convergence, leading to a substantial reduction in the total number of iterations, computations, and overall communication required for convergence.
@@
To the best of our knowledge, this is the first work to reduce the total communication cost by enabling fixed multiple communication rounds within a single iteration. Our results open a promising direction for decentralized optimization by revealing a new trade-off between the number of per-iteration communication rounds and overall convergence speed.
@@ 1.1 Related Work
Decentralized optimization has gained significant attention in large-scale machine learning, particularly in scenarios where data is distributed across multiple agents or devices without a central coordinator.
A common strategy in these methods is the use of stochastic or deterministic mixing matrices to perform local averaging of variables across neighboring nodes, facilitating global consensus through communication. Early methods such as Decentralized Gradient Descent (DGD) (DGD1, ; DGD2, ; DGDConv, ) directly extend classical gradient-based algorithms to networked environments, laying the groundwork for later developments. However, DGD suffers from slow convergence and sensitivity to step size. To address these shortcomings, a line of gradient-tracking based algorithms has been developed, including EXTRA and PG-EXTRA (EXTRA, ; PG-EXTRA, ), NIDS (NIDS, ), SONATA (SONATA, ), and (GT1, ), which incorporate correction terms to estimate the average gradient across the network.
Several of these algorithms, including (PG-EXTRA, ; NIDS, ; SONATA, ) are designed for decentralized composite optimization problems with smooth-nonsmooth structure by embedding proximal gradient steps. And the unifying analysis in (unify, ) shows that a broad family of methods—including (PG-EXTRA, ; NIDS, ; GT1, )—can be understood through a single theoretical framework and established linear convergence under strong convexity assumptions.
In addition to improving per-iteration convergence rates, many recent methods incorporate acceleration techniques such as Nesterov acceleration, leading to further improvements in both computation and communication complexity including (Multi1, ; Multi2, ; DNGD, ; ProxMudag, ).
In parallel, a different family of methods focuses on dual-based formulations. These include (Dual1, ; Dual2, ; Dual3, ) and decentralized adaptations of the Alternating Direction Method of Multipliers (ADMM) (DADMM1, ; DADMM, ; DADMM3, ; DADMM2, ) which reformulate the problem with consensus constraints and alternate between primal and dual updates. A notable development in this line is (DPADMM, ), which construct constraints based on mixing matrix into an ADMM framework for decentralized composite optimization.
ADMM and Symmetric ADMM
@@
The Alternating Direction Method of Multipliers (ADMM) is a powerful algorithmic framework for solving linearly constrained convex optimization problems with separable objective structures, and it does so without requiring smoothness assumptions. This characteristic makes ADMM particularly well-suited for composite optimization tasks involving non-smooth loss functions and regularizers. We refer readers to (ADMMEckstein, ; Boyd, ; Glowinski2014, ) for comprehensive overviews of the classical ADMM framework, and to (PADMM, ; nonergo, ; ADMMconv1, ; ADMMconv, ) for detailed convergence analyses of ADMM and its various extensions.
To improve convergence speed and numerical performance, Symmetric ADMM (S-ADMM) and its generalizations have been proposed in recent years (SADMM, ; GSADMM, ). These methods modify the standard ADMM iteration by introducing a symmetric primal-dual update structure, typically involving an additional intermediate update of the dual variable. This symmetric design allows for more balanced update dynamics between the primal and dual variables and often leads to improved practical performance. Convergence analyses for S-ADMM and its extensions have been established in works such as (GSADMMConv, ; PRSM, ).
Multi-Communication in Decentralized Optimization
@@
Incorporating multiple communication rounds per iteration has been explored in decentralized optimization for different purposes, using either fixed or adaptive strategies. Early work by (DGDT, ) showed that fixed multi-communication inherits DGD’s convergence issues and incur high communication costs, while adaptive strategies—where communication rounds increase periodically—can achieve exact convergence, albeit requiring tuning or prior knowledge.
Later methods (Multi1, ; ProxMudag, ; Multi2, ) adopted fixed multi-communication to attain optimal theoretical communication complexity. However, empirical results in (ProxMudag, ) indicate that fixed multi-round schemes are unable to reduce total communication. This limitation motivates algorithmic designs that embed multi-communication within the problem structure, rather than treating it as an external enhancement.
1.2 Notation
We denote $\mathbf{1}_{m}\in\mathbb{R}^{m}$ as the vector of all ones, and $I_{m}$ as the $m\times m$ identity matrix. The nullspace and range space of a matrix $A$ are denoted by $\mathrm{null}(A)$ and $\mathrm{span}(A)$, respectively. For any symmetric, positive definite matrix $M$ and any vector $x$ of compatible dimension, we define the weighted norm as $\|x\|_{M}:=\sqrt{\langle Mx,x\rangle}$. The distance from a point $x$ to a set $S$ with respect to $M$ is defined as $\mathrm{dist}_{M}(x,S):=\inf\left\{\left\|x-s\right\|_{M}\,\middle|\,s\in S\right\},$ and we omit the subscript $M$ when $M=I$.
The proximal operator of a convex function $g(\cdot)$ with parameter $\lambda>0$ is defined as
$\text{Prox}_{\lambda g}(v)=\underset{x\in\mathbb{R}^{m}}{\arg\min}\left(g(x)+\frac{1}{2\lambda}\|x-v\|^{2}\right).$ (1)
We use it for the proximal operator of many common functions have explicit forms.
@@ 2.1 Problem Setup
In this paper, we consider a network of $n$ agents collaboratively solving a decentralized composite optimization problem of the form:
$\min_{x\in\mathbb{R}^{d}}F(x)=\sum_{i=1}^{n}\left[f_{i}(x)+g_{i}(x)\right],$ (2)
where $f_{i}$ is a convex loss function and $g_{i}$ is a convex local regularizer both privately held by agent $i$. In cases involving a global regularizer $g$, we decompose it across the agents via $g_{i}=\lambda_{i}g$, ensuring that $\sum_{i=1}^{n}\lambda_{i}=1$. This framework models many decentralized machine learning tasks. Typical examples of loss functions include least squares, quantile loss, Huber loss, and hinge loss, while examples for regularization include the $\ell_{1}$-norm, $\ell_{2}$-norm, and elastic net.
@@ 2.2 Graph Topology
The communication structure among agents is modeled by an undirected and connected graph $G=(V,E)$, where $V=\{1,2,\dots,n\}$ is the set of agents, and an edge $(i,j)\in E$ indicates a direct communication link between agents $i$ and $j$.
The graph structure is encoded by a mixing matrix $W\in\mathbb{R}^{n\times n}$, where $W_{ij}\in[0,1]$ denotes the communication weight between agents $i$ and $j$. The mixing matrix satisfies the following assumptions:
@@
(1) $W$ is symmetric; (2) $W$ is doubly stochastic as $W\mathbf{1}=\mathbf{1}$, where $\mathbf{1}$ denotes the all-ones vector; (3) $W_{ij}>0,i\neq j$ if and only if $(i,j)\in E$ and $W_{ii}>0$ for all $i\in V$.
The assumption of the graph and corresponding mixing matrices leads to the following properties which can be derived from the Perron-Frobenius Theorem:
@@
(1) The eigenvalues of the mixing matrix satisfy $1=\lambda_{1}(W)>\lambda_{2}(W)\geq\dots\geq\lambda_{n}(W)>-1$, and the spectral gap $\rho=1-\max\{|\lambda_{2}(W)|,|\lambda_{n}(W)|\}>0$; (2) $\mathrm{null}(I_{n}-W)=\mathrm{span}\{\mathbf{1}_{m}\}$ .
See Appendix A for an example of mixing matrix based on Metropolis-Hastings weight (Weight, ).
@@ 3 Method
This section is of the design of our communication-efficient decentralized algorithm. We begin by reformulating the consensus constraint to accommodate multi-round communication within each iteration. We then describe how proximal linearization ensures decentralized updates, outline the communication strategy that minimizes per-iteration transmission, and finally present the full decentralized version of our proposed algorithm.
3.1 Reformulating Consensus Constraints
Which is crucial in applying Symmetric ADMM in decentralized optimization is to enforce agreement among all agents’ local variables. However, unlike in Distributed ADMM (Boyd, ) where a global coordinator can explicitly manage consistency, the decentralized setting requires a different strategy to ensure consensus, particularly to form suitable linear constraints. The inspiration was from the null space property of the matrix $I_{n}-W$. Define $\tilde{W}=W\otimes I_{d}$, where $\otimes$ denotes the Kronecker product. Let $u=(u_{1}^{\top},\dots,u_{n}^{\top})^{\top}\in\mathbb{R}^{nd}$ be the stacked vector of local variables across the network. It is straight to verify the following proposition since a positive spectral gap exist:
Proposition 2.
The consensus condition $u_{1}=u_{2}=\dots=u_{n}$ is equivalent to $(I_{nd}-\tilde{W}^{\top}\tilde{W})u=0$.
To integrate this consensus structure into our algorithm, we introduce the auxiliary constraint variable $v=(v_{1}^{\top},\dots,v_{n}^{\top})^{\top}\in\mathbb{R}^{nd}$ and propose the following symmetric formulation of the consensus constraint: $u=\tilde{W}v$, $\tilde{W}u=v$. This system embeds consensus into a linear structure that is both symmetric and decentralized, making it ideal for Symmetric ADMM application.
Denote $f(u)=\sum_{i=1}^{n}f_{i}(u_{i})$ and $g(v)=\sum_{i=1}^{n}g_{i}(v_{i})$, using these constraints, problem (2) can be equivalently reformulated as the following form:
$\underset{u,v\in\mathbb{R}^{nd}}{\min}f(u)+g(v)\quad s.t.\quad Au-Bv=0$ (3)
@@ 3.1 Reformulating Consensus Constraints
The ADMM algorithm proposed in (DPADMM, ) adopts a similar constraint formulation, with $A=((I_{nd}-\widetilde{W})^{1/2},I_{nd})^{\top}$ and $B=(0,I_{nd})^{\top}$. This design enables a single communication round per iteration by absorbing part of the Lagrange multiplier update. However, this asymmetry prevents the application of Symmetric ADMM, which requires a balanced primal-dual formulation.
3.2 Proximal Linearization
[…]
This choice ensures that the augmented Lagrangian terms become separable across agents, enabling fully decentralized computation. It also aligns with prior work in proximal ADMM (PADMM, ), now adapted for use in a decentralized setting. As a result, the global update form (15) can be derived and presented in Appendix B for brevity.
Given the global update rules and the symmetry of the mixing matrix $W$, we can decompose the updates into per-agent computations and derive a fully decentralized algorithm.
3.3 Communication Strategy
For convenience, we define local aggregations via neighbor communication as $\tilde{w}_{1i}^{(t)}=\sum_{j=1}^{n}W_{ji}w_{1j}^{(t)}$, $\tilde{w}_{2i}^{(t)}=\sum_{j=1}^{n}W_{ji}w_{2j}^{(t)}$, $\tilde{v}_{i}^{(t)}=\sum_{j=1}^{n}W_{ji}v_{j}^{(t)}$, $\tilde{u}_{i}^{(t)}=\sum_{j=1}^{n}W_{ji}u_{j}^{(t)}$.
The use of the quadratic form $\widetilde{W}^{\top}\widetilde{W}$ in the global update rule (15) indicates that each agent’s update involves 2-distance neighbor information. This observation leads to the following result:
Proposition 3.
For our proposed algorithm, a single iteration which contains updates to both primal variables and two updates to the Lagrange multiplier requires at least two rounds of communication to transmit the necessary information.
Given the presence of four independent $d$-dimensional variables per agent, it is essential to optimize communication scheduling, both in timing and in which variables are transmitted. This leads to two guiding principles: (1) restrict communication to two rounds per iteration, and (2) minimize transmitted data per round.
Each Symmetric ADMM iteration follows the sequence: $u$-update → first $\lambda$-update → $v$-update → second $\lambda$-update. The first update of $w_{1}$ depends on mixed $u$ information, requiring a communication step between the $u$-update and the first $w_{1}$-update. Likewise, a second communication is needed after the $v$-update to support the $w_{2}$-update. These two rounds are essential and cannot be collapsed.
To reduce communication overhead, we carefully select what is transmitted. In the first round, sending updated $u$ is sufficient for updating $w_{1}$, but not for updating $v$, which depends on more refined information. Rather than transmitting primal variables directly, we send dual variables that implicitly encode the needed content. Specifically, transmitting $a_{i}^{(t)}=w_{2i}^{(t+\frac{1}{2})}+\frac{1}{r}(w_{2i}^{(t+\frac{1}{2})}-w_{2i}^{(t)})$ allows correct computation of $v_{i}^{(t+1)}$. This positions the first communication round between the first $w_{2}$-update and the $w_{1}$-update. The second round similarly transmits $b_{i}^{(t+1)}=w_{1i}^{(t+1)}+\frac{1}{s}(w_{1i}^{(t+1)}-w_{1i}^{(t+\frac{1}{2})}).$ between the second $w_{1}$-update and the $w_{2}$-update.
Together, these insights determine the optimal placement and content of communication in each iteration, with two rounds of communication per iteration and two $d$-dimensional variables transmitted per round.

### v2

@@ Abstract
This paper focuses on decentralized composite optimization over networks without a central coordinator. We propose a novel decentralized symmetric ADMM algorithm that incorporates multiple communication rounds within each iteration, derived from a new constraint formulation that enables information exchange beyond immediate neighbors. While increasing per-iteration communication, our approach significantly reduces the total number of iterations and overall communication cost. We further design optimal communication rules that minimize the number of rounds and variables transmitted per iteration. The proposed algorithms are shown to achieve linear convergence under standard and relatively weak assumptions (e.g., metric subregularity). Extensive experiments on regression and classification tasks validate the theoretical results and demonstrate superior performance compared to existing decentralized optimization methods.
@@ 1 Introduction
In decentralized optimization, a central challenge lies in reducing the time cost of both local computation and inter-node communication. While existing algorithms differ in their local update strategies, most follow a common structural pattern: each iteration is followed by a single round of communication. This convention has seldom been challenged, primarily due to the concern that introducing multiple communication rounds per iteration would increase the overall communicational cost. Prior attempts to incorporate fixed multiple communication rounds, such as those gradient-based methods in (Jakovetić et al., 2014; Berahas et al., 2019; Ye et al., 2023; Li et al., 2018), achieve this by multi-consensus—repeatedly mixing local variables through communication. However, these methods have not demonstrated practical reductions in the total number of communication rounds. As a result, the potential for achieving a net communication reduction through non-adaptive multi-communication algorithms remains largely unexplored.
@@
In this paper, we explore incorporating multiple communication rounds within a single iteration in an approach orthogonal to existing multi-consensus schemes inspired by this observation. Rather than directly applying multiple mixing steps, we develop our algorithms by introducing linear constraints tailored for ADMM, which naturally embed multi-round communication into each iteration. To further enhance performance, we adopt a Symmetric ADMM (He et al., 2016) framework to accelerate convergence. Although this design increases the per-iteration communication cost, it enables a significantly faster convergence, leading to a substantial reduction in the total number of iterations, computations, and overall communication required for convergence.
@@
To the best of our knowledge, such a symmetric-ADMM construction with a fixed multi-round schedule and the associated theory has not been explored in prior decentralized optimization literature, and it successfully reduces the overall communication cost. Our results open a promising direction for decentralized optimization by revealing a new trade-off between the number of per-iteration communication rounds and overall convergence speed.
@@ 1.1 Related Work
Decentralized optimization has gained significant attention in large-scale machine learning, particularly in scenarios where data is distributed across multiple agents or devices without a central coordinator. A common strategy in these methods is the use of stochastic or deterministic mixing matrices to perform local averaging of variables across neighboring nodes, facilitating global consensus through communication. Early methods such as Decentralized Gradient Descent (DGD) (Nedic and Ozdaglar, 2009; Jakovetić et al., 2012; Yuan et al., 2016) directly extend classical gradient-based algorithms to networked environments, laying the groundwork for later developments. However, DGD suffers from slow convergence and sensitivity to step size. To address these shortcomings, a line of gradient-tracking based algorithms has been developed, including EXTRA and PG-EXTRA (Shi et al., 2015b; a), NIDS (Li et al., 2019), SONATA (Sun et al., 2022), and (Qu and Li, 2018), which incorporate correction terms to estimate the average gradient across the network. Several of these algorithms, including (Shi et al., 2015a; Li et al., 2019; Sun et al., 2022) are designed for decentralized composite optimization problems with smooth-nonsmooth structure by embedding proximal gradient steps. And the unifying analysis in (Xu et al., 2021) shows that a broad family of methods—including (Shi et al., 2015a; Li et al., 2019; Qu and Li, 2018)—can be understood through a single theoretical framework and established linear convergence under strong convexity assumptions. In addition to improving per-iteration convergence rates, many recent methods incorporate acceleration techniques such as Nesterov acceleration, leading to further improvements in both computation and communication complexity including (Jakovetić et al., 2014; Li et al., 2018; Qu and Li, 2020; Ye et al., 2023). In parallel, a different family of methods focuses on dual-based formulations. These include (Uribe et al., 2020; Scaman et al., 2017; Lan et al., 2017) and decentralized adaptations of the Alternating Direction Method of Multipliers (ADMM) (Wei and Ozdaglar, 2013; Shi et al., 2014; Chang et al., 2015; Aybat et al., 2018) which reformulate the problem with consensus constraints and alternate between primal and dual updates. A notable development in this line is (Wang et al., 2018), which construct constraints based on mixing matrix into an ADMM framework for decentralized composite optimization.
ADMM and Symmetric ADMM.
@@
The Alternating Direction Method of Multipliers (ADMM) is a powerful algorithmic framework for solving linearly constrained convex optimization problems with separable objective structures, and it does so without requiring smoothness assumptions. This characteristic makes ADMM particularly well-suited for composite optimization tasks involving non-smooth loss functions and regularizers. We refer readers to (Eckstein and Yao, 2015; Boyd et al., 2011; Glowinski, 2014) for comprehensive overviews of the classical ADMM framework, and to (Deng and Yin, 2016; He and Yuan, 2015; Han et al., 2018; Yuan et al., 2020) for detailed convergence analyses of ADMM and its various extensions. To improve convergence speed and numerical performance, Symmetric ADMM (S-ADMM) and its generalizations have been proposed in recent years (He et al., 2016; Bai et al., 2018). These methods modify the standard ADMM iteration by introducing a symmetric primal-dual update structure, typically involving an additional intermediate update of the dual variable. This symmetric design allows for more balanced update dynamics between the primal and dual variables and often leads to improved practical performance. Convergence analyses for S-ADMM and its extensions have been established in works such as (Bai et al., 2019; Gu et al., 2015).
Multi-Communication in Decentralized Optimization.
@@
Incorporating multiple communication rounds per iteration has been explored in decentralized optimization for different purposes, using either fixed or adaptive strategies. Early work by (Berahas et al., 2019) showed that fixed multi-communication inherits DGD’s convergence issues and incur high communication costs, while adaptive strategies—where communication rounds increase periodically—can achieve exact convergence, albeit requiring tuning or prior knowledge. Later methods (Jakovetić et al., 2014; Ye et al., 2023; Li et al., 2018) adopted fixed multi-communication to attain optimal theoretical communication complexity. However, empirical results in (Ye et al., 2023) indicate that fixed multi-round schemes are unable to reduce total communication. This limitation motivates algorithmic designs that embed multi-communication within the problem structure, rather than treating it as an external enhancement.
@@ 2.1 Problem Setup
We consider a network of $n$ agents collaboratively solving the decentralized composite optimization problem
$\min_{x\in\mathbb{R}^{d}}F(x)=\sum_{i=1}^{n}\big[f_{i}(x)+g_{i}(x)\big],$ (1)
where $f_{i}$ is a convex local loss function and $g_{i}$ is a convex local regularizer, both privately held by agent $i$. In settings involving a global regularizer $g$, we distribute it across the agents as $g_{i}=p_{i}g$ with nonnegative weights satisfying $\sum_{i=1}^{n}p_{i}=1$. This formulation captures a broad range of decentralized machine learning problems. Typical examples of loss functions include least squares, quantile loss, Huber loss, and hinge loss, while common regularizers include the $\ell_{1}$-norm, $\ell_{2}$-norm, and elastic net. Moreover, we denote the proximal operator of function $f$ as
$\text{Prox}_{\eta f}(y)=\underset{x\in\mathbb{R}^{m}}{\arg\min}\left(f(x)+\frac{1}{2\eta}\|x-y\|^{2}\right)$ (2)
Many commonly used loss and regularization functions admit closed-form proximal operators, which we leverage in the algorithmic development that follows.
@@ 2.2 Graph Topology
We model the communication pattern among agents using an undirected and connected graph $G=(V,E)$, where $V=\{1,1,\dots,n\}$ denotes the set of agents and an edge $(i,j)\in E$ represents a direct bidirectional communication link between agents $i$ and $j$. The graph structure is encoded by a mixing matrix $W\in\mathbb{R}^{n\times n}$, where $W_{ij}\in[0,1]$ specifies the communication weight assigned to agent $j$ by agent $i$. Throughout the paper, we impose the following standard assumptions on $W$:
@@
The mixing matrix $W$ satisfies: (1) $W$ is symmetric; (2) $W$ is doubly stochastic, i.e., $W\mathbf{1}=\mathbf{1}$, where $\mathbf{1}$ denotes the all-ones vector; (3) $W_{ij}>0$ for $i\neq j$ if and and only if $(i,j)\in E$, and $W_{ii}>0$ for all $i\in V$.
An example of such a matrix based on the Metropolis–Hastings weights  (Hastings, 1970) is provided in Appendix A. Under Assumption 1, the graph and its mixing matrix enjoy several well-known properties that follow from the Perron–Frobenius theorem:
@@
The following properties hold: (1) The eigenvalues of $W$ satisfy $1=\lambda_{1}(W)>\lambda_{2}(W)\geq\dots\geq\lambda_{n}(W)>-1,$ and the spectral gap $\rho=1-\max\{|\lambda_{2}(W)|,|\lambda_{n}(W)|\}$ is strictly positive; (2) The null space of $I_{n}-W$ is given by $\mathrm{null}(I_{n}-W)=\mathrm{span}\{\mathbf{1}\},$ which characterizes the consensus subspace.
@@ 3 Method
This section presents the design of our communication-efficient decentralized algorithm. We begin by reformulating the consensus constraint to enable multiple rounds of communication within each iteration. We then explain how proximal linearization facilitates decentralized updates, describe our communication scheme that minimizes per-iteration transmission, and finally introduce the full decentralized form of the proposed algorithm.
3.1 Symmetric Consensus Constraints
A key requirement for applying symmetric ADMM (He et al., 2016) in decentralized optimization is to enforce agreement among agents’ local variables. Unlike distributed ADMM (Boyd et al., 2011), where a central coordinator explicitly enforces consistency, the decentralized setting lacks a global authority and therefore requires a different mechanism for encoding consensus within linear constraints. Our approach is motivated by the spectral properties of the mixing matrix $W$, in particular the structure of the null space of $I_{n}-W$. Let $\widetilde{W}=W\otimes I_{d}$, and let $u=(u_{1}^{\top},\ldots,u_{n}^{\top})^{\top}\in\mathbb{R}^{nd}$ denote the stacked vector of local variables. Using the fact that $W$ is symmetric, doubly stochastic, and has a positive spectral gap, one can show:
Proposition 2 (Symmetric consensus constraint).
[…]
This pair of constraints in the above Proposition 2 embeds consensus into a linear, fully decentralized structure. Let $f(u)=\sum_{i=1}^{n}f_{i}(u_{i})$ and $g(v)=\sum_{i=1}^{n}g_{i}(v_{i})$. Using the above constraints, problem equation 1 is equivalently reformulated as
$\min_{u,v\in\mathbb{R}^{nd}}f(u)+g(v)\qquad\text{s.t.}\qquad Au-Bv=0,$ (3)
@@ 3.1 Symmetric Consensus Constraints
The ADMM variant proposed in (Wang et al., 2018) uses a related but asymmetric constraint formulation with $A=\big((I_{nd}-\widetilde{W})^{1/2},\,I_{nd}\big)^{\top}$ and $B=(0,\,I_{nd})^{\top}$. This design enables a single communication round per iteration by absorbing part of the dual update. However, the inherent asymmetry prevents the use of symmetric ADMM, which requires a balanced primal–dual structure.
3.2 Graph-Aware Proximal Linearization
Let $\lambda=(\lambda_{1},\lambda_{2})$ denote the Lagrange multipliers associated with the two equality constraints in equation 3, let $\beta>0$ be the augmented Lagrangian penalty parameter, and let $r,s>0$ be the symmetric dual update steps. The generalized symmetric ADMM (see, e.g., He et al., 2016; Bai et al., 2018; Deng and Yin, 2016) applied to the reformulated problem equation 3 takes the form
$\left\{\begin{aligned} u^{(t+1)}&=\arg\min_{u\in\mathbb{R}^{nd}}\;f(u)-\langle\lambda^{(t)},Au\rangle+\frac{\beta}{2}\|Au-Bv^{(t)}\|^{2}+\frac{1}{2}\|u-u^{(t)}\|_{Q}^{2},\\ \lambda^{(t+\frac{1}{2})}&=\lambda^{(t)}-r\beta\big(Au^{(t+1)}-Bv^{(t)}\big),\\ v^{(t+1)}&=\arg\min_{v\in\mathbb{R}^{nd}}\;g(v)+\langle\lambda^{(t+\frac{1}{2})},Bv\rangle+\frac{\beta}{2}\|Au^{(t+1)}-Bv\|^{2}+\frac{1}{2}\|v-v^{(t)}\|_{Q}^{2},\\ \lambda^{(t+1)}&=\lambda^{(t+\frac{1}{2})}-s\beta\big(Au^{(t+1)}-Bv^{(t+1)}\big).\end{aligned}\right$. (4)
To efficiently linearize the subproblems and remove the quadratic coupling introduced by the terms $\|Au-Bv^{(t)}\|^{2}$ and $\|Au^{(t+1)}-Bv\|^{2}$, we introduce a graph-aware positive definite proximal matrix
$Q\;=\;\beta\big((1+\tau)I_{nd}-\widetilde{W}^{\top}\widetilde{W}\big),\qquad\tau>0.$
This choice ensures that the augmented Lagrangian becomes separable across agents, enabling fully decentralized updates of both primal variables. Then, by expanding the symmetric ADMM updates in equation 4, the updates can be rewritten in the following decomposable form:
$\left\{\begin{aligned} u^{(t+1)}&=\underset{{u\in\mathbb{R}^{nd}}}{\arg\min}\,f(u)-\Big\langle\widetilde{W}\lambda_{1}^{(t)}+\lambda_{2}^{(t)}+\beta\!\left(2\widetilde{W}v^{(t)}+(1+\tau)u^{(t)}-\widetilde{W}^{\!\top}\widetilde{W}u^{(t)}\right),\,u\Big\rangle\\ &\quad\;\;+\beta\!\left(1+\frac{\tau}{2}\right)\!\|u\|^{2},\\[6.0pt] \lambda_{1}^{(t+\frac{1}{2})}&=\lambda_{1}^{(t)}-r\beta\big(\widetilde{W}u^{(t+1)}-v^{(t)}\big),\;\lambda_{2}^{(t+\frac{1}{2})}=\lambda_{2}^{(t)}-r\beta\big(u^{(t+1)}-\widetilde{W}v^{(t)}\big),\\[6.0pt] v^{(t+1)}&=\underset{v\in\mathbb{R}^{nd}}{\arg\min}\,g(v)-\Big\langle\beta\!\left(2\widetilde{W}u^{(t+1)}+(1+\tau)v^{(t)}-\widetilde{W}^{\!\top}\widetilde{W}v^{(t)}\right)-\lambda_{1}^{(t+\frac{1}{2})}-\widetilde{W}\lambda_{2}^{(t+\frac{1}{2})},v\Big\rangle\\ &+\beta\!\left(1+\frac{\tau}{2}\right)\!\|v\|^{2},\\[6.0pt] \lambda_{1}^{(t+1)}&=\lambda_{1}^{(t+\frac{1}{2})}-s\beta\big(\widetilde{W}u^{(t+1)}-v^{(t+1)}\big),\;\lambda_{2}^{(t+1)}=\lambda_{2}^{(t+\frac{1}{2})}-s\beta\big(u^{(t+1)}-\widetilde{W}v^{(t+1)}\big).\end{aligned}\right$. (5)
This graph-aware linearization eliminates quadratic coupling, restores decomposability, and yields update rules that are fully decentralized and structurally balanced across the two primal blocks. For a detailed derivation of the general update form, we refer the reader to Appendix B.
3.3 Optimal Double-Communication Structure
We now describe how to implement the decomposable symmetric ADMM updates in equation 5 in a fully decentralized manner. There are four blocks of variables $u,\lambda_{1},\lambda_{2},v\in\mathbb{R}^{nd}$. We write
$u=(u_{1}^{\top},\dots,u_{n}^{\top})^{\top},v=(v_{1}^{\top},\dots,v_{n}^{\top})^{\top},\lambda_{1}=(\lambda_{1,1}^{\top},\dots,\lambda_{1,n}^{\top})^{\top},\lambda_{2}=(\lambda_{2,1}^{\top},\dots,\lambda_{2,n}^{\top})^{\top}$
with $u_{i},v_{i},\lambda_{1,i},\lambda_{2,i}\in\mathbb{R}^{d}$ stored locally at agent $i$, and communication is restricted to exchanging weighted neighbor-averaged values determined by the mixing matrix $W$. For convenience, for any stacked variable $z=(z_{1}^{\top},\dots,z_{n}^{\top})^{\top}$ we use the notation $\tilde{z}_{i}^{(t)}\;=\;\sum_{j\in\mathcal{N}_{i}}W_{ij}z_{j}^{(t)}.$
A key structural feature of the symmetric ADMM updates is that the quadratic term $\widetilde{W}^{\top}\widetilde{W}$ in equation 5 requires 2-hop neighbor information. Consequently, one full iteration necessarily involves two distinct rounds of communication, and the communication pattern must therefore be scheduled with care. Our design follows two ordered principles: (1) achieve the minimal number of rounds by restricting communication to exactly two rounds per iteration; (2) within each round, minimize the amount of transmitted data.
The first principle is enforced because $\lambda_{1}^{(t+\frac{1}{2})}$ requires aggregated $u^{(t+1)}$, and $\lambda_{2}^{(t+1)}$ requires aggregated $v^{(t+1)}$, fixing the two communication rounds between these updates. Following this structure, we organize each iteration into two update groups—Group 1 and Group 2—separated by two communication rounds. To satisfy the second principle, we transmit carefully chosen dual-variable combinations rather than raw primal variables, reducing the communication in each round to only two $d$-dimensional vectors and achieving the minimal cost compatible with the symmetric ADMM structure. The overall communication–computation workflow is described as follows:
Group 1 update and Communication 1.
At iteration $t$, agent $i$ begins with its current local variables $(u_{i}^{(t)},v_{i}^{(t)},\lambda_{1,i}^{(t)},\lambda_{2,i}^{(t-\frac{1}{2})})$ and the cached mixed values $\tilde{v}_{i}^{(t)}$ and $\tilde{b}_{i}^{(t)}$ from the previous iteration. In Group 1, agent $i$ first updates $\lambda_{2,i}^{(t)}$, then computes $u_{i}^{(t+1)}$ via the proximal operator of $f_{i}$, and finally forms the intermediate dual iterate $\lambda_{2,i}^{(t+\frac{1}{2})}$.
At this point, the first communication round occurs. Instead of transmitting all primal variables, each agent constructs the auxiliary message $a_{i}^{(t+1)}=\lambda_{2,i}^{(t+\frac{1}{2})}+\frac{1}{r}\bigl(\lambda_{2,i}^{(t+\frac{1}{2})}-\lambda_{2,i}^{(t)}\bigr),$ and broadcasts the pair $(u_{i}^{(t+1)},a_{i}^{(t+1)})$ to its neighbors. The received messages are aggregated to provide all mixed information needed for the Group 2 updates.
Group 2 update and Communication 2.
Using $\tilde{u}_{i}^{(t+1)}$ and $\tilde{a}_{i}^{(t+1)}$, agent $i$ performs the Group 2 update: it updates $\lambda_{1,i}^{(t+\frac{1}{2})}$, computes $v_{i}^{(t+1)}$ via the proximal operator of $g_{i}$, and then forms $\lambda_{1,i}^{(t+1)}$.
In the second communication round, each agent constructs $b_{i}^{(t+1)}=2\lambda_{1,i}^{(t+1)}-\lambda_{1,i}^{(t+\frac{1}{2})},$ and broadcasts $(v_{i}^{(t+1)},b_{i}^{(t+1)})$ to its neighbors, and received messages become inputs for the next Group 1 update.